OpenStudy (anonymous):

Find the period, range, and amplitude of the cosine function. y = 3/2 cos t/2

OpenStudy (anonymous):

Can you post the choices?

OpenStudy (anonymous):

Amplitude = 3/2 Period = 4 * pi Range = [ -3/2, 3/2]

OpenStudy (anonymous):

period = 4pi, range: –three-halves ≤ y ≤three-halves; amplitude = three-halves period = one-half, range: –three-halves ≤ y ≤ three-halves; amplitude = –three-halves period = 4pi, range: –three-halves ≤ y ≤three-halves; amplitude = –three-halves period = one-half, range: y ≤ three-halves; amplitude = three-halves

OpenStudy (anonymous):

A